One number is 4 times another.The sum of the numbers is 65. Find the numbers.
Answer by venugopalramana(3286)

One number is 4 times another.The sum of the numbers is 65. Find the numbers.
let one number=x
four times that number =4x=the other number
so sum of the 2 numbers=x+4x=65
5x=65
x=65/5=13
so one number is 13 and the other number is 4x=4*13=52